# Spinal open books and symplectic fillings with exotic fibers

### v1

@@ 4.2. Uniqueness of the local model
The local model describes a foliation such that in the asymptotic coordinates near each puncture {±$\{\pm 1\}\in\mathbb{C}$, the curves have leading terms being complex multiples of $e^{-it}$.
@@
To understand the asymptotic presentation of the curves $u_{c}$ near the punctures in the local model in Definition 4.2, we will perform a change of coordinates on the domain from {±$\mathbb{C}\setminus\{\pm 1\}$ given by $z^{\prime}:=z-1$. This will give asymptotic coordinates near the puncture at 1, and it will be symmetric for the -1 case. We will also write the image of $u_{c}$ in ×$\mathbb{R}\times S^{1}\times\mathbb{C}$ coordinates as in Lemma 4.1. The Reeb orbit in question, corresponding to $\{0\}\times\mathbb{C}$ in our previous coordinates or {}×$\{\infty\}\times S^{1}\times\{0\}$ in the current coordinates, has period 1 and in the model, the level $3$-manifold $\{s\}\times S^{1}\times\mathbb{C}$ has the standard metric, hence we can use the trivial connection, hence the exponential map is just identity.
@@
As we are using the trivial connection in {.}×$\{.\}\times S^{1}\times\mathbb{R}^{2}$, the exponential map is identity under identification of $\xi_{\theta}$ with {.}×$\{.\}\times\{\theta\}\times\mathbb{R}^{2}$. So we can work with the above expression to understand the leading terms in the asymptotic expansion. The ×$\mathbb{R}\times S^{1}$ coordinate for the curve is $u_{c,\mathbb{R}\times S^{1}}=(-\ln(|cz^{\prime}(z^{\prime}+2)|),-\phi_{1})$. Let us call that $(s,t)$. We then focus on the $\mathbb{R}^{2}$ coordinates, call it $u_{c,\mathbb{R}^{2}}$. We have that
@@
Now we prove uniqueness of such a local model for an exotic fiber. To prepare for the proof for the uniqueness of the local model, we first establish the some notations and Lemma 4.8 detailing a few asymptotic properties of a regular fiber in the neighborhood of an exotic fiber.
Recall that ^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})$ is the specific moduli space of pseudoholomorphic curves in $(\widehat{W},\widehat{\omega},\widehat{J})$ described in §3.2. In particular, it allows a partition into regular, singular and exotic moduli spaces together with ^$\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+})$ and there is a continuous surjection
@@
by Theorem 3.2. To establish notation, given a neighborhood $D\cong\mathbb{C}$ of an exotic fiber in the moduli space ^$\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+})$, we let $\widehat{u_{c}}:=\Pi^{-1}(c)$ where $c\in\mathbb{C}$ denote the pseudoholomorphic map from $\Sigma_{c}$ into $\widehat{W}$ corresponding to either a regular fiber when $c\neq 0$ and an exotic fiber when $c=0$.
For $c\neq 0$, consider now a twice-punctured disk $P\subset\Sigma_{c}$ for which there is an almost complex chart to {±$\mathbb{C}\setminus\{\pm 1\}$. In addition, identify the neighborhood $\widehat{N}\subset\widehat{W}$ of the elliptic orbit $\gamma_{e}$ with ($(\mathbb{C}\setminus\{0\})\times\mathbb{C}$ using Lemma 4.1. The image of $\partial P$ under $\widehat{u_{c}}$, by continuity, naturally lands in the closure $\overline{\widehat{N}}$. Note that this closure can be understood as a product of $\mathbb{C}\setminus\{0\}$ and $\mathbb{C}$ both compactified at infinity, and the boundary of $\overline{\widehat{N}}$ belongs to the interior of $\widehat{W}$, away from the infinite end. In particular, the projection maps $\Pi_{0}$ and $\Pi$ on $\widehat{N}$ make sense in its closure.
@@
Consider an exotic fiber in a completed filling $(\widehat{W},\widehat{\omega},\widehat{J})$. Let $\widehat{N}$ be as above – a noncompact neighborhood of the doubly covered orbit $\gamma_{e}$ in the exotic fiber. Then there exists a smooth coordinate chart $\phi:\widehat{N}\to\mathbb{C}\setminus\{0\}\times\mathbb{C}$ such that the intersection of the the nearby neighborhood regular fibers with $\widehat{N}$ is sent to the foliation $\0\mathcal{F}_{0}:=\{u_{c}|c\in\mathbb{C}\backslash\{0\}\}$, whose leaves are described in (8) and are the non-zero level sets of the map
@@
Given $c\in\mathbb{C}\setminus\{0\}$, let $u_{c}:=\widehat{u_{c}}|_{P}$ be a regular curve in a neighborhood of the exotic fiber in the moduli space ^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})$, restricted to a non-compact neighborhood of the two asymptotic end at $\gamma_{e}$ as above. By the continuous surjection in Proposition 3.2, or more directly, by the SFT compactness in the proof of Proposition 3.2, $u_{c}$ can be parameterized by maps of the form
@@
in a non-compact neighborhood of the collapsing punctures, where $f_{c}$ and $p_{c}$ are both holomorphic functions on {±$\mathbb{C}\setminus\{\pm 1\}$. See Figure 6.
Now, by assumption of the asymptotic neighborhood, $q_{c}$ can be extended to have removable singularities at ±$\pm 1$ with value 00 and is non-vanishing on {±$\mathbb{C}\setminus\{\pm 1\}$. This means $q_{c}$ is a multiple of $(z^{2}-1)$, call it
@@
where $\alpha:\mathbb{C}\setminus\{0\}\to\mathbb{C}\setminus\{0\}$ is a smooth function in $c$. The $p_{c}$ coordinate is determined by the asymptotics of the holomorphic curves near the two ends at ±$\pm 1$, whereas the coordinate $q_{c}$ determines a simultaneous identification between $z\in P$ and $(s,t)$ as in Equation (2), near ±$\pm 1$, for all the curves $u_{c}$ for $c\in\mathbb{C}\setminus\{0\}$.
@@
In addition, by Lemma 4.8, we also have that $c_{N}(\widehat{u_{c}})=0$. Therefore, we may apply Lemma 3.4 to obtain a local diffeomorphism between the moduli space of regular curves in the neighborhood of an exotic fiber at the asymptotic end $\gamma$ and $\mathbb{C}\setminus\{0\}$. The local diffeomorphism allows us to pick asymptotic cylinders that we call $Z_{c}^{1}$ and $Z_{c}^{-1}$ in the neighborhood $D$ of an exotic fiber in the moduli space ^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})$, for which the leading eigenfunctions have coefficients $\sqrt{c}$ and $-\sqrt{c}$, respectively.
Step 2 (first deformation): Observe that Lemma 4.7 also shows that our local model consists of curves with leading eigenfunction coefficients ±$\pm\sqrt{c}$, where the $\sqrt{c}$ coefficient corresponds to the asymptotic end of the curve $u_{c}$ near the puncture at +1, while $-\sqrt{c}$ corresponds to the the asymptotic end of $u_{c}$ near the puncture at $-1$.
Now we modify the foliation ^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})$ smoothly on a neighborhood $N\subset\widehat{N}$ near the asymptotic ends of the curves. This can be done by a similar analysis of Lemma 5.5 in [Wen10b]. Choose $R_{1}$ sufficiently large so that the tangent spaces of the curves in $N$ is uniformly close to the tangent space of the asymptotic orbit cylinders in $[R_{1},\infty)\times Y$. Now we modify the rest of the asymptotic terms of $Z_{c}^{1}$ other than the leading term by cutoff functions to match the expansion in Lemma 4.7, for the curve with leading eigenfunction coefficient $\sqrt{c}$. As in Lemma 5.5 in [Wen10b], the modified foliation is still $J^{\prime}$-holomorphic, for some $J^{\prime}$ uniformly closed to $\widehat{J}$, so it’s also tamed by $\widehat{\omega}$. Let $\omega_{\epsilon}$ be a symplectic form tamed by $J_{\epsilon}$. We then have that the modification in the pseudoholomorphic foliation is induced by a symplectic deformation $\omega_{t}=(1-t)\widehat{\omega}+t\omega_{\epsilon}$.
Simultaneously, for $Z_{c}^{-1}$, we modify the rest of the asymptotic terms to match the expansion in Lemma 4.7, for the curve with leading eigenfunction coefficient $-\sqrt{c}$. Doing this for all $c$ means that near ±$\pm 1$, the curves $u_{c}$ agree with the curves in Lemma 4.7. By analytic continuation, it follows that they agree over the whole domain.
@@
Parameterize $\widehat{N}$ by $\0(r,\theta,p,q)\in\mathbb{R}\times S^{1}\times\mathbb{R}^{2}\cong(\mathbb{C}\backslash\{0\})\times\mathbb{C}$ by the identification via Lemma 4.1. In particular, as $r\to 0$, $t\to\infty$. Choose $\epsilon>0$ small enough so that $\widehat{N}_{r\leq\epsilon}$ corresponds to the symplectization in the double completion $\widehat{E}$ as described in §3.2, where all the curves in the foliation are cylinders over their asymptotic ends. Since $\widehat{N}$ is the symplectization over a neighborhood of a Reeb orbit, by the discussion in [LHMW20, §3], $\widehat{N}_{r\leq\epsilon}\subset(-1,\infty)\times M_{\Sigma}\subset\widehat{E}$. Comparing the symplectic structures and stable Hamiltonian structures in the local model $\widehat{N}$ and in $(-1,\infty)\times M_{\Sigma}\subset\widehat{E}$, it follows that the sets $\{r_{0},\theta_{0}\}\times\mathbb{C}$ can be identified with the vertebrae in $\{r_{0}\}\times M_{\Sigma}$. Therefore, the intersection of the vertebrae with $\widehat{N}$ are holomorphic, and the projection map $\Pi:\widehat{W}\to\mathcal{M}$ takes the form $(z_{1},z_{2})\mapsto z_{2}^{2}-z_{1}$.
We now fix a representative $\Sigma$ of the vertebra, whose intersection with $\widehat{N}$ is $\Sigma_{0}:=\{r=r_{0},\theta=0\}$ where $r_{0}<\epsilon$. We now examine the intersections of $\Sigma_{0}$ with the fibers $\operatorname{\text{im}}(u_{c})$. Parameterizing {±$\mathbb{C}\setminus\{\pm 1\}$ by $(x,y)$, with the natural identification $z=x+iy$, we see that
@@
[…]
@@ 2. II)
Case (ii): Now, on $\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})$, consider the subset $W_{\operatorname{\text{exot}}}$. Every cylindrical end of the foliation that does not intersect $W_{\operatorname{\text{exot}}}$ covers their asymptotic Reeb orbit once, by the same argument as the above paragraph. Thus $\Pi|_{\Sigma_{t}}$ is again an honest covering map restricted to $\Sigma_{t}\bigcap\bigg{(}\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})\setminus W_{\operatorname{\text{exot}}}\bigg{)}$. Since every curve in the foliation that lives in $\bigg{(}\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})\bigg{)}$ has exactly $(k-2)$ cylindrical ends in $\bigg{(}\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})\setminus W_{\operatorname{\text{exot}}}\bigg{)}$, this is a $(k-2)$-degree covering map.
@@ 5. Construction of strong fillings
According to [EVV10, Lemma 3.2], the page $P$ can be made convex such that the dividing set consists of boundary parallel arcs. We can isotope $\gamma$ on $P$ to avoid the dividing set. Then, by applying the Legendrian realization principle, we can assume $\gamma$ is Legendrian with the contact framing equal to the framing coming from $P$. Since a positive Dehn twist about $\gamma$ corresponds to a $(-1)$-surgery with respect to the surface framing, it corresponds to a Legendrian surgery. Therefore, a Weinstein 2-handle attachment corresponding to the Legendrian surgery produces a Weinstein cobordism from $(M,\xi)$ to $(M^{\prime},\xi^{\prime})$. ∎
@@
By Theorem 2.8, a bordered Lefschetz fibration without a vanishing cycle admits a Stein structure. According to Hayden [Hay21], the boundary of a positive multisection in a Stein surface is quasipositive. Therefore, we can apply Proposition 5.5 and the 4-manifold constructed in Step 1 admits a symplectic structure which is an exact filling of the boundary. The 2-handles attached along vanishing cycles admits a Weinstein structure by Lemma 5.6. A Weinstein cobordism preserves the exact fillability and this establishes Theorem 5.2.
@@
[GR79, p.130][Sim63, Theorem 1] If $X$ is a normal Stein space of complex dimension two and $H$ is an analytic hypersurface in $X$, then $X\backslash H$ is Stein.
@@
[BO01] Any symplectic surface in $(\mathbb{D}^{4},\omega_{st})$ is symplectically isotopic to a positive multisection in the Lefschetz fibration $\mathbb{D}^{2}\times\mathbb{D}^{2}\to\mathbb{D}^{2}$, and it is further symplectically isotopic to a complex curve.
@@
By Lemma 5.3 and its proof, $(M,\xi)$ has the spinal open book induced by a filling $E$ by PANLF which is constructed by the following. First, consider the complement of a multisection $S$ in $\natural_{n}S^{1}\times\mathbb{D}^{3}$, with boundary a quasipositive link $K$ with respect to a planar open book supporting (#$(\#_{n}S^{1}\times S^{2},\xi_{st})$, for some $n$. Here $S$ is determined by $\{\Sigma_{i}\}_{n_{i}\neq 0}$, since we know $B=\mathbb{D}^{2}$. Second, we add $2$-handles along vanishing cycles corresponding to $\phi$ to obtain $E$.
Now in order to see that this procedure in fact gives rise to a Stein filling, we need to examine the multisection complement $E_{1}:=\natural_{n}S^{1}\times\mathbb{D}^{3}\setminus S$ in the first step. Observe that $E_{1}$ can also be obtained by removing a multisection $S^{\prime}\cup\sqcup_{n}D^{2}$ from the $\mathbb{D}^{2}\times\mathbb{D}^{2}$ fibration of $\mathbb{D}^{4}$, where $S^{\prime}$ is homeomorphic to $S$. The boundary of $S^{\prime}\cup\sqcup_{n}\mathbb{D}^{2}$ is a quasipositive link $K^{\prime}$ in the disk open book obtained by considering the braid $K$ in the planar open book supporting (#$(\#_{n}S^{1}\times S^{2},\xi_{st})$, and the boundaries of the disks correspond to all but one boundary components in the same planar open book – these give trivial components in the link $K^{\prime}$.
By Observation 5.4 and Theorem 5.8, we know that $S^{\prime}\cup\sqcup_{n}\mathbb{D}^{2}$ is isotopic to a complex curve in $\mathbb{D}^{4}$, which we denote as $\mathcal{S}$. Applying Theorem 5.7, we obtain that $E_{1}$, the complement of $\mathcal{S}$ in $\mathbb{D}^{4}$, is Stein. Then, as $E$ is obtained by attaching Weinstein handles to $E_{1}$, corresponding to vanishing cycles, $E$ is Stein as well. Note that by the proof of Lemma 1 in [Sim63], we may pick the new plurisubharmonic function given in Theorem 5.7 to have a level set whose contact structure agrees with the boundary of the complement of the neighborhood of $\mathcal{S}$ in $\mathbb{D}^{4}$, which by construction is the contact structure supported by the given spinal open book decomposition that we desire. ∎

### v2

@@ 4.2. Uniqueness of the local model
The local model describes a foliation such that in the asymptotic coordinates near each puncture $\{\pm 1\}\in\mathbb{C}$, the curves have leading terms being complex multiples of $e^{-it}$.
@@
To understand the asymptotic presentation of the curves $u_{c}$ near the punctures in the local model in Definition 4.2, we will perform a change of coordinates on the domain from $\mathbb{C}\setminus\{\pm 1\}$ given by $z^{\prime}:=z-1$. This will give asymptotic coordinates near the puncture at 1, and it will be symmetric for the -1 case. We will also write the image of $u_{c}$ in $\mathbb{R}\times S^{1}\times\mathbb{C}$ coordinates as in Lemma 4.1. The Reeb orbit in question, corresponding to $\{0\}\times\mathbb{C}$ in our previous coordinates or $\{\infty\}\times S^{1}\times\{0\}$ in the current coordinates, has period 1 and in the model, the level $3$-manifold $\{s\}\times S^{1}\times\mathbb{C}$ has the standard metric, hence we can use the trivial connection, hence the exponential map is just identity.
@@
As we are using the trivial connection in $\{.\}\times S^{1}\times\mathbb{R}^{2}$, the exponential map is identity under identification of $\xi_{\theta}$ with $\{.\}\times\{\theta\}\times\mathbb{R}^{2}$. So we can work with the above expression to understand the leading terms in the asymptotic expansion. The $\mathbb{R}\times S^{1}$ coordinate for the curve is $u_{c,\mathbb{R}\times S^{1}}=(-\ln(|cz^{\prime}(z^{\prime}+2)|),-\phi_{1})$. Let us call that $(s,t)$. We then focus on the $\mathbb{R}^{2}$ coordinates, call it $u_{c,\mathbb{R}^{2}}$. We have that
@@
Now we prove uniqueness of such a local model for an exotic fiber. To prepare for the proof for the uniqueness of the local model, we first establish some notations and Lemma 4.8 detailing a few asymptotic properties of a regular fiber in the neighborhood of an exotic fiber.
[…]
@@
by Theorem 3.2. To establish notation, given a neighborhood $D\cong\mathbb{C}$ of an exotic fiber in the moduli space $\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+})$, we let $\widehat{u_{c}}:=\Pi^{-1}(c)$ where $c\in\mathbb{C}$ denote the pseudoholomorphic map from $\Sigma_{c}$ into $\widehat{W}$ corresponding to either a regular fiber when $c\neq 0$ and an exotic fiber when $c=0$.
For $c\neq 0$, consider now a twice-punctured disk $P\subset\Sigma_{c}$ for which there is an almost complex chart to $\mathbb{C}\setminus\{\pm 1\}$. In addition, identify the neighborhood $\widehat{N}\subset\widehat{W}$ of the elliptic orbit $\gamma_{e}$ with $(\mathbb{C}\setminus\{0\})\times\mathbb{C}$ using Lemma 4.1. The image of $\partial P$ under $\widehat{u_{c}}$, by continuity, naturally lands in the closure $\overline{\widehat{N}}$. Note that this closure can be understood as a product of $\mathbb{C}\setminus\{0\}$ and $\mathbb{C}$ both compactified at infinity, and the boundary of $\overline{\widehat{N}}$ belongs to the interior of $\widehat{W}$, away from the infinite end. In particular, the projection maps $\Pi_{0}$ and $\Pi$ on $\widehat{N}$ make sense in its closure.
@@
[…]
@@
[…]
@@
in a non-compact neighborhood of the collapsing punctures, where $f_{c}$ and $p_{c}$ are both holomorphic functions on $\mathbb{C}\setminus\{\pm 1\}$. See Figure 6.
Now, by assumption of the asymptotic neighborhood, $q_{c}$ can be extended to have removable singularities at $\pm 1$ with value 0 and is non-vanishing on $\mathbb{C}\setminus\{\pm 1\}$. This means $q_{c}$ is a multiple of $(z^{2}-1)$, call it
@@
where $\alpha:\mathbb{C}\setminus\{0\}\to\mathbb{C}\setminus\{0\}$ is a smooth function in $c$. The $p_{c}$ coordinate is determined by the asymptotics of the holomorphic curves near the two ends at $\pm 1$, whereas the coordinate $q_{c}$ determines a simultaneous identification between $z\in P$ and $(s,t)$ as in Equation (2), near $\pm 1$, for all the curves $u_{c}$ for $c\in\mathbb{C}\setminus\{0\}$.
@@
[…]
Step 2 (first deformation): Observe that Lemma 4.7 also shows that our local model consists of curves with leading eigenfunction coefficients $\pm\sqrt{c}$, where the $\sqrt{c}$ coefficient corresponds to the asymptotic end of the curve $u_{c}$ near the puncture at +1, while $\sqrt{c}$ corresponds to the the asymptotic end of $u_{c}$ near the puncture at 1-1.
[…]
Simultaneously, for $Z_{c}^{-1}$, we modify the rest of the asymptotic terms to match the expansion in Lemma 4.7, for the curve with leading eigenfunction coefficient $\sqrt{c}$. Doing this for all $c$ means that near $\pm 1$, the curves $u_{c}$ agree with the curves in Lemma 4.7. By analytic continuation, it follows that they agree over the whole domain.
@@
We now fix a representative $\Sigma$ of the vertebra, whose intersection with $\widehat{N}$ is $\Sigma_{0}:=\{r=r_{0},\theta=0\}$ where $r_{0}<\epsilon$. We now examine the intersections of $\Sigma_{0}$ with the fibers $\operatorname{\text{im}}(u_{c})$. Parameterizing $\mathbb{C}\setminus\{\pm 1\}$ by $(x,y)$, with the natural identification $z=x+iy$, we see that
$u_{c}=\begin{pmatrix}\operatorname{\text{Re}}(\sqrt{c})x-\operatorname{\text{Im}}(\sqrt{c})y\\ \operatorname{\text{Re}}(\sqrt{c})y+\operatorname{\text{Im}}(\sqrt{c})x\\ \operatorname{\text{Re}}(c)(x^{2}-y^{2}-1)-\operatorname{\text{Im}}(c)2xy\\ \operatorname{\text{Re}}(c)2xy+\operatorname{\text{Im}}(c)(x^{2}-y^{2}-1)\end{pmatrix}$
@@
$Du_{c}|_{(x,y)}=\begin{pmatrix}\operatorname{\text{Re}}(\sqrt{c})&-\operatorname{\text{Im}}(\sqrt{c})\\ \operatorname{\text{Im}}(\sqrt{c})&\operatorname{\text{Re}}(\sqrt{c})\\ 2x\operatorname{\text{Re}}(c)-2y\operatorname{\text{Im}}(c)&-2y\operatorname{\text{Re}}(c)-2x\operatorname{\text{Im}}(c)\\ 2y\operatorname{\text{Re}}(c)+2x\operatorname{\text{Im}}(c)&2x\operatorname{\text{Re}}(c)-2y\operatorname{\text{Im}}(c)\end{pmatrix}$.
@@ 2. II)
Case (ii): Now, on $\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})$, consider the subset $W_{\operatorname{\text{exot}}}$. Every cylindrical end of the foliation that does not intersect $W_{\operatorname{\text{exot}}}$ covers their asymptotic Reeb orbit once, by the same argument as the above paragraph. Thus $\Pi|_{\Sigma_{t}}$ is again an honest covering map restricted to $\Sigma_{t}\bigcap\bigg{(}\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})\setminus W_{\operatorname{\text{exot}}}\bigg{)}$. Since every curve in the foliation that lives in $\bigg{(}\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})\bigg{)}$ has exactly (k2)(k-2) cylindrical ends in $\bigg{(}\Pi|_{\Sigma_{t}}^{-1}(D_{\operatorname{\text{exot}}})\setminus W_{\operatorname{\text{exot}}}\bigg{)}$, this is a (k2)(k-2)-degree covering map.
@@ 5. Construction of strong fillings
According to [EVV10, Lemma 3.2], the page $P$ can be made convex such that the dividing set consists of boundary parallel arcs. We can isotope $\gamma$ on $P$ to avoid the dividing set. Then, by applying the Legendrian realization principle, we can assume $\gamma$ is Legendrian with the contact framing equal to the framing coming from $P$. Since a positive Dehn twist about $\gamma$ corresponds to a (1)(-1)-surgery with respect to the surface framing, it corresponds to a Legendrian surgery. Therefore, a Weinstein 2-handle attachment corresponding to the Legendrian surgery produces a Weinstein cobordism from $(M,\xi)$ to $(M^{\prime},\xi^{\prime})$. ∎
@@
By Theorem 2.8, a bordered Lefschetz fibration without a vanishing cycle admits a Stein structure. According to Hayden [Hay21], the boundary of a positive multisection in a Stein surface is a quasipositive braid. Therefore, we can apply Proposition 5.5 and the 4-manifold constructed in Step 1 admits a symplectic structure which is an exact filling of the boundary. The 2-handles attached along vanishing cycles admits a Weinstein structure by Lemma 5.6. A Weinstein cobordism preserves the exact fillability and this establishes Theorem 5.2.
@@
[…]
Notice that in the trivial Lefschetz fibration $D^{4}=D^{2}\times D^{2}$, a positive multisection of the base $D^{2}$ is equivalent to the quasipositive surface; see [BO01, Definition 1.1]. Thus we can restate [BO01, Theorem 2] as follows:
@@
[BO01] Any symplectic surface in $(\mathbb{D}^{4},\omega_{st})$ is symplectically isotopic to a positive multisection in the Lefschetz fibration $\mathbb{D}^{2}\times\mathbb{D}^{2}\to\mathbb{D}^{2}$.
@@
By Lemma 5.3 and its proof, $(M,\xi)$ has the spinal open book induced by a filling $E$ by PANLF which is constructed by the following. First, consider the complement of a multisection $S$ in $\natural_{n}S^{1}\times\mathbb{D}^{3}$, with boundary a quasipositive link $K$ with respect to a planar open book supporting $(\#_{n}S^{1}\times S^{2},\xi_{st})$, for some $n$. Here $S$ is determined by $\{\Sigma_{i}\}_{n_{i}\neq 0}$, since we know $B=\mathbb{D}^{2}$. Second, we add $2$-handles along vanishing cycles corresponding to $\phi$ to obtain $E$.
Now in order to see that this procedure in fact gives rise to a Stein filling, we need to examine the multisection complement $E_{1}:=\natural_{n}S^{1}\times\mathbb{D}^{3}\setminus S$ in the first step. Observe that $E_{1}$ can also be obtained by removing a symplectic multisection $S^{\prime}\cup\sqcup_{n}D^{2}$ from the $\mathbb{D}^{2}\times\mathbb{D}^{2}$ fibration of $\mathbb{D}^{4}$, where $S^{\prime}$ is homeomorphic to $S$. The boundary of $S^{\prime}\cup\sqcup_{n}\mathbb{D}^{2}$ is a quasipositive link $K^{\prime}$ in the disk open book obtained by considering the braid $K$ in the planar open book supporting $(\#_{n}S^{1}\times S^{2},\xi_{st})$, and the boundaries of the disks correspond to all but one boundary components in the same planar open book – these give trivial components in the link $K^{\prime}$.
By Theorem 5.8, we know that $S^{\prime}\cup\sqcup_{n}\mathbb{D}^{2}$ is symplectically isotopic to a positive multisection in $\mathbb{D}^{2}\times\mathbb{D}^{2}$, denoted as $\mathcal{S^{\prime}}$. Furthermore, by [Rud83] we know that $\mathcal{S}^{\prime}$ is smoothly isotopic to a complex curve, denoted as $\mathcal{S}$. Applying Theorem 5.7, we obtain that $E_{1}$, the complement of $\mathcal{S}$ in $\mathbb{D}^{4}$, is Stein. In fact, we may pick an explicit plurisubharmonic function on $\mathbb{D}^{2}\times\mathbb{D}^{2}\setminus\mathcal{S}$ to be $f:=\max(-\log(d_{\mathcal{S}}),|z_{1}|^{2}+|z_{2}|^{2})$ where $d_{\mathcal{S}}$ denotes the distance function to the set $\mathcal{S}$ and $z_{1},z_{2}$ denote the two complex coordinates on $\mathbb{D}^{2}\times\mathbb{D}^{2}$.
We now check that a level set of $f$ is compatible with a contact structure supported by the spinal open book obtained by removing $\mathcal{S}$. Notice that the induced contact form on a regular level set of $f$ is $\lambda:=-df\circ J_{0}$ where $J_{0}$ is the standard complex structure on $\mathbb{D}^{4}$. We first see that $d\lambda$ is a symplectic form on each regular fiber and hence each page of the open book. Second, we claim that $\lambda$ evaluates positively on each fiber of the spine. Notice that removing a small neighborhood of $\mathcal{S}$ introduces some boundary components on each fiber surface $\{p\}\times\mathbb{D}^{2}$ for $p\in\mathbb{D}^{2}$. The $S^{1}$-fibration on the spine coincides with the boundaries of the fibers away from the exotic curves, and near the exotic curves can be made arbitrarily close to the boundaries of the fiber. Thus to show the claim, it is sufficient to show that $\lambda$ evaluates positively on tangent vectors of each boundary component of each fiber surface. Let $X$ be a non-vanishing tangent vector field along a boundary component of a fiber surface, inducing the boundary orientation. Then $J_{0}(X)$ points away from the the multisection $\mathcal{S}$, so the function $f$ decreases along the integral curves of $J_{0}(X)$. Therefore, it follows that $\lambda(X)>0$ and this completes the proof that $\ker(\lambda)$ is supported by the induced spinal open book on the boundary.
Finally, as $E$ is obtained by attaching Weinstein handles to $E_{1}$, corresponding to vanishing cycles, $E$ is Stein as well and the boundary contact structure is supported by the induced spinal open book. ∎